# A novel language model for predicting serious adverse event results in clinical trials from their prospective registrations

### v1

@@ Abstract
Objectives: With accurate estimates of expected safety results, clinical trials could be designed to avoid terminations and limit exposing participants to unnecessary risks. We evaluated methods for predicting serious adverse event (SAE) results in clinical trials using information only from their registrations prior to the trial.
Material and Methods: We analysed 22,107 two-arm parallel interventional clinical trials from ClinicalTrials.gov with structured summary results. Two prediction models were developed: a classifier predicting will experimental arm have higher SAE rates (area under the receiver operating characteristic curve; AUC) than control arm, and a regression model to predict the proportion of SAEs in control arms (root mean squared error; RMSE). A transfer learning approach using pretrained language models (e.g., ClinicalT5, BioBERT) was used for feature extraction, combined with downstream model for prediction. To maintain semantic representation in long trial texts exceeding localised language model input limits, a sliding window method was developed for embedding extraction.
Results: The best model (ClinicalT5+Transformer+MLP) had 77.6% AUC predicting which trial arm has a higher proportion of patients with SAEs. When predicting proportion of participants experiencing SAE in the control arm, the same model achieved RMSE of 18.6%. The sliding window approach consistently outperformed methods without it. Across 12 classifiers, the average absolute AUC increase was 2.00%; across 12 regressors, the average absolute RMSE reduction was 1.58%.
Discussion: Summary results data available at ClinicalTrials.gov remains underutilised. The potential to estimate results of trials before they start is an opportunity to improve trial design and flag discrepancies between expected and reported safety results.
@@ 1 Background
Clinical trials are designed to establish evidence of the efficacy and safety of new treatments and track its effectiveness and monitor its safety post-approval[1]. Clinical trials are required to monitor participants for adverse events2. Serious adverse events (SAE) are defined as unexpected medical occurrences that can lead to death, hospitalization, or significant medical intervention[2]. Serious adverse events (SAE) are defined as unexpected medical occurrences that can lead to death, hospitalization, or significant medical intervention[2].
@@
Taking advantage of the semi-structured nature of clinical trial registration data, there has been broad application of machine learning on the data. Prior work using ClinicalTrials.gov as the basis for training machine learning models to predict trial termination[10], publication likelihood[11], new therapeutics progressing through trial phases towards approval[12], and trial durations for specific diseases[13]. Despite the importance of clinical trials and their registrations for establishing evidence of safety, we know of no studies that have sought to predict safety results of trials from registrations.
The possibility of predicting safety results of clinical trials from their prospective registrations suggests a range of novel use cases. Prediction of safety results could be used to assess the design of trials before they start or flag trials for further investigation where reported safety results vary substantially from what was expected. Where safety results can be accurately modelled by learning from similar study arms in completed trials, there may also be potential to construct synthetic control arms[14].
Our aim was to predict safety results about SAEs for clinical trials from their prospective trial registrations. The specific aims were to predict will the experimental arm have a higher proportion of patients experiencing SAEs than the control arm and estimate the proportion of participants who will experience SAEs in study control arms.
@@ 2.1 Dataset and preprocessing
From a set of 292,133 interventional studies on ClinicalTrials.gov, 21% (61,330 trials) had summary results reported on ClinicalTrials.gov (as of March 5, 2024). We included interventional studies that were parallel group clinical trials where participants were randomised to one of two groups (called trial arms), which included a set of 22,107 trials (Figure 1). Clinical trials with this design typically compare a new or experimental treatment against a previous standard of care or a placebo.
@@
For the classification task of predicting whether the experimental arm will have a higher proportion of participants experiencing SAEs compared to the control arm, we used downsampling to address data imbalance. We randomly downsample the majority class (class 0), resulting in 11,540 trials (5,770 trials in each class). For the regression task of predicting the proportion of participants experiencing a SAE in the control arm, we observed a skewed distribution with an overrepresentation of low-proportion trials. We grouped trials into ten bins based on SAE proportion and randomly sampled up to 1,000 trials per bin, yielding a final set of 5,216 trials.
@@ 2.1.1 Model architecture
The model architecture was based on a transfer learning approach. Following preprocessing steps for each trial, its embedding was extracted using pretrained language model. We tested a range of language models (see Table 1). The language model weights were frozen during training to preserve the semantic knowledge learned from large-scale corpora and to reduce computational cost. The generated embedding was then fed into a downstream prediction model, which was used for the task-specific training by calculating loss and updating weights accordingly. The model then produced a prediction based on the processed features.
@@
BGE-m3 [21] 8192 A language model trained on diverse text data, representing models not specialised in the biomedical domain.
@@
For the downstream prediction model, we evaluated three different modal architecture for both classification and regression tasks: K-Nearest Neighbours (KNN), Multi-layer Perceptron (MLP), and Transformer Encoder combined with MLP. In the KNN model, predictions were based on the majority label among the k-nearest neighbours in the feature space using cosine similarity (with k = 20 for regression and k = 60 for classification, selected via grid search). The MLP model was a feedforward neural network consisting of 12 fully connected layers with non-linear activation functions. The Transformer Encoder + MLP model included a Transformer Encoder with 12 layers and 8 attention heads to capture contextual relationships within the input embedding, followed by a 3-layer MLP for final prediction. The hyperparameters (k for KNN, number of layers for MLP, and layers for Transformer Encoder + MLP) were tuned using a validation set through grid search over various configurations.
@@
Localised language models performance can degrade when processing lengthy input due to their context length limitation[22] Some methods have been developed to improve how language models handle long contexts, such as changing model architecture or positional encoding strategies[23, 24, 25, 26]. While these strategies offer improvements, they need substantial additional training or fine-tuning of the language model[27].
@@
Our feature extraction process employs a sliding window approach to ensure complete document coverage despite model context limitations. It involves splitting the input text into overlapping chunks based on the model’s maximum context length, then processing each chunk individually, and combining them to produce a final embedding (Figure 2). This allows us to process the long registration document while avoiding the need for extensive retraining of the language model. For instance, with BioBERT’s 512-token limit, we used a 256-token stride, allowing half of the content from each chunk to carry over into the next one. We selected 50% overlap to balance computational efficiency with context preservation, though this parameter could be optimised in future work. To establish a baseline for our experiments, we adopted a common embedding extraction approach used in previous studies involving extracting embedding on electronic health record text and clinical trial registration text[11, 28]. In the common embedding approach, the entire input text of a single trial is processed as one continuous sequence.
@@
Figure 2: In a comparison example of the baseline approach and the sliding window approach, the sliding window approach a window size 2 and stride 1. The sliding window example showed above has a window size $2$ and stride $1$. $T_{i}$: The $i$-th token after tokenising the input text. $n$: Total number of tokens generated from the input text. $h$: Number of units in the last hidden state of the language model (e.g., 768 for BERT). $w$: Total number of chunks generated by the sliding window. ${e_{i}}^{k}$: The $i$-th element of the embedding from chunk $k$. $E_{i}$: The $i$-th element in the final embedding.
@@ 2.1.2 Evaluation and outcome measures
For model training, we set the learning rate to 0.0001 for all our models. To optimise the training process, we used the AdamW optimizer[29]. We used the Cross Entropy Loss function during training for classification task and Mean Square Error Loss function for regression task training. Hyperparameters for the model architecture were chosen based on grid search through various configuration. Each classification model was trained for 20 epochs, and each regression model was trained for 40 epochs, all chosen based on the elbow point observed during training. To ensure reproducibility, we used a random seed of 42 for train and test set split, classification train set downsampling, and test set bootstrap resampling.
Accuracy, F1 Score and AUC (Area Under the Receiver Operating Characteristic Curve) were used as measures of performance in the classification task. For the regression task, performance was measured using mean absolute error (MAE; the average absolute difference between the predicted and actual values) and root mean square (RMSE; the square root of the average squared differences between the predicted and actual values), which penalises larger errors more heavily.
@@ 3 Results
The best performing configuration for the regression task was the same as the classification task (Table 3). This configuration achieved a RMSE of 18.56%. For completeness, MAE was also reported (14.27%), but we primarily evaluated performance based on RMSE, given that it penalises larger errors more heavily and is more sensitive to clinically meaningful prediction deviations. Statistical testing using bootstrapped Wilcoxon signed-rank tests comparing both metrics indicated a significant difference in performance between all configurations (Supplementary Material Table 2).
@@ 4 Discussion
In this study we demonstrated the feasibility of predicting aspects of results of clinical trials related to SAEs, using only information from the prospective registration of the trial. This included a novel sliding window method to support embedding extraction, which outperformed a baseline approach consistently across most configurations. We also found that domain pretrained models, especially those trained on biomedical or clinical texts, consistently outperformed general models in processing clinical registration data.
Other studies have used ClinicalTrials.gov data to predict trial characteristics or outcomes. Previous work has focused on predicting trial publication likelihood (64.1% F1 score)[11], predicting which trials will progress to later phase trials (67.2-74.3% accuracy)[12], trial termination (66.3% accuracy)[10], and phase 1 trial duration for specific diseases (74% accuracy)[13]. In contrast, our work is the first to predict SAE results, achieving 77.6% AUC for identifying whether the experimental arm will have higher rates of SAEs compared to the control arm, and 18.6% RMSE for predicting the proportion of participants experiencing SAEs in the control arm. These results demonstrate the feasibility of predicting safety outcomes from registration data alone.
@@
The sliding window approach combined with transfer learning is relatively simple and provides a practical solution for researchers working with clinical trial registration text, and may be suitable for other tasks where tasks need to make sense of longer documents. We found that the solution is particularly well-suited to clinical trial registration text, as these documents are organised into distinct sections (eligibility criteria, intervention details, outcomes, sponsor information, etc.). Unlike narrative text where understanding might require connecting information across the entire document, each section in trial registrations tends to be relatively self-contained.
The implications of this work include helping trial investigators identify potential issues with the trial design before patient enrolment. For example, when predictions indicate high SAE rates in control arms, investigators may consider alternative approaches such as synthetic or historical control arms, potentially reducing unnecessary patient exposure to harmful treatments[14, 30] Our approach could also assist regulatory bodies with oversight. For example, they might use such predictions before trial approval to identify high-risk trials that may require enhanced monitoring of reporting[31, 32], or compare expected results to results that are finally reported in ClinicalTrials.gov or published articles, looking for unusual discrepancies as a potential indicator of problematic reporting.
@@
During label construction for our classification task, we also applied the Chi-square test to determine statistical significance of SAE differences. We labelled trials into three classes: non-significant, or if significant, as having higher SAEs in either the control or experimental arm. However, this process resulted in a highly imbalanced dataset, with most trials classified as having non-significant differences. Statistical significance depends on multiple factors including things like the sample size which are often limited in clinical trials due to cost constraints, leading to limited statistical power for detecting differences. The imbalanced dataset limited the suitability for model fine-tuning and our proposed training approach. Future work could explore methods to address the class imbalance and develop specialized approaches for the multi-class task that enable models to better distinguish trials with statistically significant differences in SAE proportions between study arms.
@@ 6.1 Detailed information on data preprocessing
  • Categorising Trial Arms. Each trial arm was labelled as either “experimental” or “control” using regular expressions based on trial title and description. This approach allowed us to efficiently identify the control and experimental arms from a large number of trials.
To validate our trial arm categorisation, we manually reviewed the five trials, and each was correctly identified for control and experimental arms.
@@
6.2 Detailed P-values for classification and regression task
@@
Note: Identical p-values arise because the Wilcoxon signed-rank test yields the same outcome whenever the signs and relative ranks of differences are consistent across comparisons. Our test has sample size 30, and the significance level is set at 0.01.

### v2

@@ Abstract
Objectives: With accurate estimates of expected safety results, clinical trials could be better designed and monitored. We evaluated methods for predicting serious adverse event (SAE) results in clinical trials using information only from their registrations prior to the trial.
Material and Methods: We analyzed 22,107 two-arm parallel interventional clinical trials from ClinicalTrials.gov with structured summary results. Two prediction models were developed: a classifier predicting whether a greater proportion of participants in an experimental arm would have SAEs (area under the receiver operating characteristic curve; AUC) compared to the control arm, and a regression model to predict the proportion of participants with SAEs in the control arms (root mean squared error; RMSE). A transfer learning approach using pretrained language models (e.g., ClinicalT5, BioBERT) was used for feature extraction, combined with a downstream model for prediction. To maintain semantic representation in long trial texts exceeding localized language model input limits, a sliding window method was developed for embedding extraction.
Results: The best model (ClinicalT5+Transformer+MLP) had 77.6% AUC when predicting which trial arm had a higher proportion of SAEs. When predicting SAE proportion in the control arm, the same model achieved RMSE of 18.6%. The sliding window approach consistently outperformed direct comparisons. Across 12 classifiers, the average absolute AUC increase was 2.00%, and absolute RMSE reduction was 1.58% across 12 regressors.
Discussion: Summary results data from ClinicalTrials.gov remains underutilized. Predicted results of publicly reported trials provides an opportunity to identify discrepancies between expected and reported safety results.
@@ 1 Background
Clinical trials are designed to establish evidence of the efficacy and safety of new treatments and track their effectiveness and monitor their safety post-approval[1]. Clinical trials are required to monitor participants for adverse events and report these results[2], but data about adverse events are underutilized even where summary results are made publicly available.
@@
Taking advantage of the semi-structured nature of clinical trial registration data, there has been broad application of machine learning on the data. Prior work using ClinicalTrials.gov as the basis for training machine learning models to predict trial termination[10], publication likelihood[11], new therapeutics progressing through trial phases towards approval[12], and trial durations for specific diseases[13]. Despite the importance of clinical trials and their registrations for establishing evidence of safety, we know of no studies that have sought to predict safety related results of trials from registrations, such as serious adverse events.
Predictive models of SAE results for clinical trials may be of value for a range of novel use cases. For example, by predicting the expected SAE results of a trial, it may be possible to flag trials for further investigation where reported SAE results vary substantially from what was expected. If other safety results can be accurately modeled by learning from similar study arms in completed trials, future opportunities may include the construction of synthetic control arms[14].
Our aim was to predict SAEs for clinical trials from their prospective trial registrations. The specific aims were to predict will the experimental arm have a higher proportion of patients experiencing SAEs than the control arm and estimate the proportion of participants who will experience SAEs in study control arms.
@@ 2.1 Dataset and preprocessing
From a set of 292,133 interventional studies on ClinicalTrials.gov, 21% (61,330 trials) had summary results reported on ClinicalTrials.gov (as of March 5, 2024). We included interventional studies that were parallel group clinical trials where participants were randomized to one of two groups (called trial arms), which included a set of 22,107 trials (Figure 1). Clinical trials with this design typically compare a new or experimental treatment against a previous standard of care or a placebo.
@@
We also extracted factors from structured data that might be associated with differences in SAE results across trial arms. We selected factors that were used in other clinical trial outcome prediction studies[10, 11, 12], and augmented these with other factors that might be expected to be associated with higher or lower proportions of SAEs in the treatment versus placebo study arm (Supplementary Material Table 1).
For the classification task of predicting whether the experimental arm will have a higher proportion of participants experiencing SAEs compared to the control arm, we used downsampling to address data imbalance. We randomly downsample the majority class (class 0), resulting in 11,542 trials (5,771 trials in each class). For the regression task of predicting the proportion of participants experiencing a SAE in the control arm, we observed a skewed distribution with an overrepresentation of low-proportion trials. We grouped trials into ten bins based on SAE proportion and randomly sampled up to 1,000 trials per bin, yielding a final set of 5,192 trials.
@@ 2.1.1 Model architecture
The model architecture was based on a transfer learning approach. Following preprocessing steps for each trial, its embedding was extracted using pretrained language model (see Table 1). The language model weights were frozen during training to preserve the semantic knowledge learned from large-scale corpora and to reduce computational cost. The generated embedding was then fed into a downstream prediction model, which was used for the task-specific training by calculating loss and updating weights accordingly. The model then produced a prediction based on the processed features.
@@
BGE-m3 [21] 8192 A language model trained on diverse text data, representing models not specialized in the biomedical domain.
@@
For the downstream prediction model, we evaluated three different modal architecture for both classification and regression tasks: K-Nearest Neighbors (KNN), Multi-layer Perceptron (MLP), and Transformer Encoder combined with MLP. In the KNN model, predictions were based on the majority label among the k-nearest neighbors in the feature space using cosine similarity (with k = 20 for regression and k = 60 for classification, selected via grid search). The MLP model was a feedforward neural network consisting of 12 fully connected layers with non-linear activation functions. The Transformer Encoder + MLP model included a Transformer Encoder with 12 layers and 8 attention heads to capture contextual relationships within the input embedding, followed by a 3-layer MLP for final prediction. The hyperparameters (k for KNN, number of layers for MLP, and layers for Transformer Encoder + MLP) were tuned using a validation set through grid search over various configurations.
@@
Localized language models performance can degrade when processing lengthy input due to their context length limitation[22]. Some methods have been developed to improve how language models handle long contexts, such as changing model architecture or positional encoding strategies[23, 24, 25, 26]. While these strategies offer improvements, they need substantial additional training or fine-tuning of the language model[27].
@@
Our feature extraction process employs a sliding window approach to ensure complete document coverage despite model context limitations. It involves splitting the input text into overlapping chunks based on the model’s maximum context length, then processing each chunk individually, and combining them to produce a final embedding (Figure 2). This allows us to process the long registration document while avoiding the need for extensive retraining of the language model. For instance, with BioBERT’s 512-token limit, we used a 256-token stride, allowing half of the content from each chunk to carry over into the next one. We selected 50% overlap to balance computational efficiency with context preservation, though this parameter could be optimized in future work. To establish a baseline for our experiments, we adopted a common embedding extraction approach used in previous studies involving extracting embedding on electronic health record text and clinical trial registration text[11, 28]. In the common embedding approach, the entire input text of a single trial is processed as one continuous sequence.
@@
Figure 2: In a comparison example of the baseline approach and the sliding window approach, the sliding window approach a window size 2 and stride 1. The sliding window example showed above has a window size 22 and stride 11. $T_{i}$: The $i$-th token after tokenizing the input text. $n$: Total number of tokens generated from the input text. $h$: Number of units in the last hidden state of the language model (e.g., 768 for BERT). $w$: Total number of chunks generated by the sliding window. ${e_{i}}^{k}$: The $i$-th element of the embedding from chunk $k$. $E_{i}$: The $i$-th element in the final embedding.
@@ 2.1.2 Evaluation and outcome measures
For model training, we set the learning rate to 0.0001 for all our models. To optimize the training process, we used the AdamW optimizer[29]. We used the Cross Entropy Loss function during training for classification task and Mean Square Error Loss function for regression task training. Hyperparameters for the model architecture were chosen based on grid search through various configuration. Each classification model was trained for 20 epochs, and each regression model was trained for 40 epochs, all chosen based on the elbow point observed during training. To ensure reproducibility, we used a random seed of 42 for train and test set split, classification train set downsampling, and test set bootstrap resampling.
Accuracy, F1 Score and AUC (Area Under the Receiver Operating Characteristic Curve) were used as measures of performance in the classification task. For the regression task, performance was measured using mean absolute error (MAE; the average absolute difference between the predicted and actual values) and root mean square (RMSE; the square root of the average squared differences between the predicted and actual values), which penalizes larger errors more heavily.
@@ 3 Results
From the 11,542 trials included in the evaluation of classification task, 33% (3,860 of 11,542) used a placebo control compared to an active comparator. Phase 3 studies (30%, 3,466 of 11,542) and Phase 2 studies (24%, 2,717 of 11,542) were most common. Most studies had funding from industry (56%, 6,475 of 11,542), most had a trial site in the United States (72%, 8,343 of 11,542), and a minority had a Data Monitoring Committee in place (44%, 5,090 of 11,542). Most factors were significantly associated with differences in SAE proportions across arms (Table 2).
Group Odds Ratio (95% CI) P-value, Fisher’s Exact Test Trials had higher proportion of SAE from the Group (%)
Trial with Placebo 1.44 (1.33, 1.55) $8.77\times 10^{-20}$ 2161 of 3860 (56%)
Trial in Phase 1 0.44 (0.36, 0.53) $2.57\times 10^{-19}$ 197 of 514 (38%)
Trial in Phase 2 1.08 (0.99, 1.19) $9.94\times 10^{-2}$ 1597 of 2717 (59%)
Trial in Phase 3 1.78 (1.63, 1.95) $3.53\times 10^{-37}$ 2276 of 3466 (66%)
Trial in Phase 4 0.53 (0.48, 0.59) $1.59\times 10^{-32}$ 799 of 1775 (45%)
Trial related to Cancer 2.70 (2.43, 2.99) $7.05\times 10^{-85}$ 1404 of 2019 (70%)
Trial hiring Female-only patients 0.76 (0.67, 0.86) $1.79\times 10^{-5}$ 472 of 1079 (44%)
Trial hiring Male-only patients 0.98 (0.79, 1.21) $8.72\times 10^{-1}$ 177 of 358 (49%)
Trial hiring Healthy volunteers 0.35 (0.32, 0.39) $9.43\times 10^{-85}$ 530 of 1814 (29%)
Trial has Data Monitoring Committee 1.67 (1.54, 1.80) $6.33\times 10^{-38}$ 2825 of 5090 (56%)
Small trial (<<100) 0.42 (0.39, 0.46) $1.03\times 10^{-113}$ 2048 of 5307 (39%)
Large trial (>>100) 2.36 (2.19, 2.55) $1.03\times 10^{-113}$ 3687 of 6172 (60%)
Trial with Industry Funding 3.04 (2.81, 3.28) $6.98\times 10^{-185}$ 4006 of 6475 (62%)
Trial Location in USA 0.94 (0.87, 1.02) $1.40\times 10^{-1}$ 4138 of 8348 (50%)
Trial Masking (Double Blind) 0.98 (0.89, 1.08) $6.82\times 10^{-1}$ 1046 of 2111 (50%)
FDA regulated 1.15 (1.05, 1.26) $2.29\times 10^{-3}$ 1320 of 2504 (53%)
Trial population Child Only 0.73 (0.64, 0.85) $3.29\times 10^{-5}$ 342 of 798 (43%)
Trial population Older Adult Only 1.15 (0.84, 1.56) $4.30\times 10^{-1}$ 87 of 163 (53%)
Table 2: Structured clinical trial information and associations with higher rates of serious adverse events in the treatment/experimental arm
For the classification task, a baseline random forest model using structured information showed reasonable performance (AUC: 69.27%, F1-score: 64.25%, Accuracy: 63.27%), predicting whether the treatment arm had a higher proportion of patients with SAEs.
@@
The best performing configuration for the regression task was the same as the classification task (Table 4). This configuration achieved a RMSE of 18.56%. For completeness, MAE was also reported (14.27%), but we primarily evaluated performance based on RMSE, given that it penalizes larger errors more heavily and is more sensitive to clinically meaningful prediction deviations. Statistical testing using bootstrapped Wilcoxon signed-rank tests comparing both metrics indicated a significant difference in performance between all configurations (Supplementary Material Table 3).
@@ 4 Discussion
In this study we demonstrated the feasibility of predicting aspects of results of clinical trials related to SAEs, using only information from the prospective registration of the trial. The predictive models that incorporated detailed text information outperformed baseline models that only used structured information. We also found that domain pretrained models, especially those trained on biomedical or clinical texts, consistently outperformed general models in processing clinical registration data.
Other studies have used ClinicalTrials.gov data to predict trial characteristics or outcomes. Previous work has focused on predicting trial publication likelihood (71.9% AUC)[11], predicting which trials will progress to later phase trials (76.0-87.3% AUC)[12], trial termination (72.6% AUC)[10], and phase 1 trial duration for specific diseases (76.8% AUC)[13]. In contrast, our work is the first to predict SAE results, achieving 77.6% AUC for identifying whether the experimental arm will have higher rates of SAEs compared to the control arm, and 18.6% RMSE for predicting the proportion of participants experiencing SAEs in the control arm. These results demonstrate the feasibility of predicting safety outcomes from registration data alone.
@@
The sliding window approach combined with transfer learning is relatively simple and provides a practical solution for researchers working with clinical trial registration text and may be suitable for other tasks where tasks need to make sense of longer documents. We found that the solution is particularly well-suited to clinical trial registration text, as these documents are organized into distinct sections (eligibility criteria, intervention details, outcomes, sponsor information, etc.). Unlike narrative text where understanding might require connecting information across the entire document, each section in trial registrations tends to be relatively self-contained.
The model we present here may be of value for regulatory oversight. For example, they might use predictions based on trial protocols prior to approval to identify high-risk trials that may require enhanced monitoring of reporting[30, 31]. Another possible use case might be to compare expected results to results that are finally reported in ClinicalTrials.gov or published articles, looking for unusual discrepancies as a flag for further investigation.
@@
During label construction for our classification task, we also applied the Chi-square test to determine statistical significance of SAE differences. We labeled trials into three classes: non-significant, or if significant, as having higher SAEs in either the control or experimental arm. However, this process resulted in a highly imbalanced dataset, with most trials classified as having non-significant differences. Statistical significance depends on multiple factors including things like the sample size which are often limited in clinical trials due to cost constraints, leading to limited statistical power for detecting differences. The imbalanced dataset limited the suitability for model fine-tuning and our proposed training approach. Future work could explore methods to address the class imbalance and develop specialized approaches for the multi-class task that enable models to better distinguish trials with statistically significant differences in SAE proportions between study arms.
@@ 6.1 Detailed information on data preprocessing
  • Categorizing Trial Arms. Each trial arm was labeled as either “experimental” or “control” using regular expressions based on trial title and description. This approach allowed us to efficiently identify the control and experimental arms from a large number of trials.
To validate our trial arm categorization, we manually reviewed the five trials, and each was correctly identified for control and experimental arms.
@@
Feature Description
trial size Total number of participants enrolled in the trial (numeric)
size category Trial size classification based on enrollment: 0=Small (<<100 participants), 1=Large ($\geq$100 participants), -1=missing
placebo Binary indicator for whether the control arm uses placebo/sham (1) or active comparator/standard care (0)
phase Clinical trial phase encoded as ordinal values: Phase 1 (1), Phase 2 (2), Phase 3 (3), Phase 4 (4), missing (-1)
cancer Binary indicator for whether the trial studies cancer or cancer-related conditions (1=yes, 0=no)
sex Participant sex eligibility encoded as: All sexes (0), Male only (1), Female only (2), missing (-1)
healthy volunteers Binary indicator for whether the trial accepts healthy volunteers (1=yes, 0=no, -1=missing)
has dmc Binary indicator for presence of a Data Monitoring Committee (1=yes, 0=no, -1=missing)
industry funded Binary indicator for industry sponsorship or collaboration (1=industry involved, 0=non-industry only)
is_usa_trial Binary indicator for whether any trial site is located in the United States (1=yes, 0=no)
num_countries Number of unique countries where the trial is conducted (numeric, 0 if missing)
double_blind Binary indicator for double-blind masking design (1=double blind, 0=other masking types or open label)
is_fda_regulated Binary indicator for FDA-regulated drug or device (1=FDA regulated, 0=not FDA regulated)
standard age Standardized age categories from trial registration (0=Child only, 1=Older adult only, 2=Rest/mixed ages, -1=missing)
Table 1: Feature used in developing the base random forest model. Missing values are encoded as -1 for all features.
6.2 Detailed P-values for classification and regression task  [^1] Identical p-values arise because the Wilcoxon signed-rank test yields the same outcome whenever the signs and relative ranks of differences are consistent across comparisons. Our test has sample size 30, and the significance level is set at 0.01.